1) Find the slope of a straight line that passes through the given pair of points:
(0,6) and (4,5)
2) Write the equation of a line that has slope = (-1/3) and passes through (3, - 4)
3) Find an equation of the line that passes through (1, - 4) and is perpendicular to
y = (-1/4)x + 4. Write answer in slope-intercept form
thank you

Answer by checkley71(8403)

slope is (y2-y1)/(x2-x1)
(5-6)/(4-0)
-1/4 is the slope
-------------------------------------------------------------
Y=mX+b
-4=-1/3*3+b
-4=-1+bB
b=-4+1
b=-3 THUS THE EQUATION OF THIS LINE IS
Y=-1/3X-3
----------------------------------------------------------------------
Y=4X+b OR
-4=4*1+b
-4=4+b
b=-8 THEN THE EQUATION OF THIS LINE & PERPENDICULAR TO ONE WITH A SLOPR OF
(-1/4) IS
Y=4X-8